OpenStudy (anonymous):

Any good with Trignometry? Please help with this problem:

OpenStudy (anonymous):

Heres the problem: \[(\sqrt[3]{x+1})^{3}-1\]

OpenStudy (anonymous):

Where's the trig?

OpenStudy (anonymous):

Its above.

OpenStudy (anonymous):

Unless I'm missing something, the answer's "x".

OpenStudy (anonymous):

haha, that is the answer. But I'm trying to figure out how to get to that answer.

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

Idk. I don't really understand how to solve it.

OpenStudy (anonymous):

Well, you're just cubing the cubed root of "x+1", which gets you back to where you started (i.e. "x+1"). And then you subtract "1", leaving you with "x".

OpenStudy (anonymous):

but ... you have an index. =\

OpenStudy (anonymous):

Well ... the root.

OpenStudy (anonymous):

Thats what killed me.

OpenStudy (anonymous):

What's confusing you?

OpenStudy (anonymous):

The root of 3.

OpenStudy (anonymous):

Oh, yeah, that just means "cubed root". To make things easier, you can always convert that to exponents.

OpenStudy (anonymous):

So, if you see: \[\sqrt[3]{x}\] You can rewrite it as: \[x ^{\frac{ 1 }{ 3 }}\] i.e. to "x" raised to the one-third power.

OpenStudy (anonymous):

Ohhh. So lets work this out together.

OpenStudy (anonymous):

\[\sqrt[3]{(3)^{3}-1}\]

OpenStudy (anonymous):

cubed root of 26

OpenStudy (anonymous):

Hahaa. WTH? How are you getting that!!!!! hahaa.

OpenStudy (anonymous):

BREAK IT DOWN ... all the way for me man!

OpenStudy (anonymous):

If you have a calculator handy you can write: \[(26)^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

How did you get 26 though.

OpenStudy (anonymous):

NVM!

OpenStudy (anonymous):

I just did it in my head. (3)^3-1 = 26

OpenStudy (anonymous):

Ok so where are we right now? Meaning: what else is a little foggy for you?

OpenStudy (anonymous):

Im going to solve this problem: \[(\sqrt[3]{x+1})^{3}-1\]

OpenStudy (anonymous):

Im guessing the answer is just X ?

OpenStudy (anonymous):

Yeah, it should be...

OpenStudy (anonymous):

Sweet! Im starting to get the hang of this now.

OpenStudy (anonymous):

Great! If you have any other questions just ask.

OpenStudy (anonymous):

I dont understand how the root comes into the answer and sometimes it cancels out. For example: \[(\sqrt[3]{x+1})^3-1 = x\] BUT : \[(\sqrt[3]{(3)^3-1} = x\] = \[(\sqrt[3]{26}\]

OpenStudy (anonymous):

Yeah--good question. It's because those two expressions on the lefthand side are NOT equal.

OpenStudy (anonymous):

Think of it this way. Actually, just forget about roots for a second. Is \[(x+1)^{3}\] the same as: \[x^{3}+1\]

OpenStudy (anonymous):

Sometimes, the two expressions MIGHT be equal but are they ALWAYS equal? No way!

OpenStudy (anonymous):

oooooh. Sweet !

OpenStudy (anonymous):

(x+1) ^3 isn't the same as x^3+1

OpenStudy (anonymous):

No, \[(x+1)^{3} = x ^{3}+3x ^{2}+3x+1\]

OpenStudy (anonymous):

That's called an IDENTITY because it is always true no matter what value you choose for "x".

OpenStudy (anonymous):

And in trig you'll see a lot of Trig IDENTITIES

OpenStudy (anonymous):

Oh okay. So that mean that when Im solving the root values, I have to watch out then. For some reason it feels like i need a little clarification on remembering when and when not to bring down the root.

OpenStudy (anonymous):

The identities thing is understandable though.